OpenStudy (anonymous):

You roll two dice. What is the probability of the event that you roll a total of two or twelve on the two dice?

OpenStudy (anonymous):

0oh, thats a good one.

OpenStudy (anonymous):

you have to work out every possible total for each type of roll, then see how many times twelve is possible, can you do it?

OpenStudy (anonymous):

2/36 or 1/18

OpenStudy (anonymous):

was that wrong, what i said, anyone else?

OpenStudy (anonymous):

what you said is right, but there is only one way to roll a 12. ditto for a 2.

OpenStudy (anonymous):

i think impusle made it two because two die are used...

OpenStudy (anonymous):

impulse made it a 2 because there are 2 ways to roll 2 or 12 with 2 dice.

OpenStudy (anonymous):

There are 36 possible rolls that are all independent of the 36, 2 of them are 2 and 12

OpenStudy (anonymous):

oh,, two possible rolls, that explains it.

myininaya (myininaya):

Possible rows are: die 1 die 2 1 1 1 2 1 3 1 4 1 5 1 6 2 1 2 2 2 3 2 4 2 5 2 6 ....and so on so there are 6+6+6+6+6+6=6(6)=36 P(rolling 2)=1/36 P(rolling 12)=1/36 P(rolling 2 or 12)=1/36+1/36=2/36=1/18

OpenStudy (anonymous):

there you go lol